# InSpire: Vision-Language-Action Models with Intrinsic Spatial Reasoning

### v2

@@ 3.1 Simulation Experiments
LIBERO-90 LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
@@
3.2 Real-world Experiments
The advantages of our InSpire approach are clearly demonstrated through simulation experiments. Here, we conduct extensive experiments on real-world tasks to assess InSpire’s practical effectiveness.
@@ 3.2 Real-world Experiments
Experimental Setup. Due to the difficulty in collecting huge pools of real-world behavioral demonstrations, it is challenging to train a model that displays generalization and robustness on varied real-world tasks, whether starting from scratch or finetuning a pretrained model lacking exposure to substantial real-world manipulation data. Hence, we employ $\pi_{0}$-FAST, the current best-performing open-source VLA, as the baseline model for real-world evaluations. We carefully design 10 seen tasks and 5 unseen tasks that focus on evaluating the VLA’s performance along multiple dimensions: spatial reasoning, interacting with novel objects and scenes, and following unknown instructions, as shown in Appendix C.2. The $\pi_{0}$-FAST model is full-finetuned using the 10 seen tasks, each consisting of training 10 training trajectories. During inference, we conduct 10 trials per task and randomize the configurations and orientations of task-specific objects for each trial. We use an AGILEX PiPER 6DOF robot arm. Other hyperparameters are consistent with those used in the simulation experiments.
Results and Analysis. Fig. 4 reports the success rates as well as time costs of $\pi_{0}$-FAST [32] integrated w/ or w/o our InSpire on 10 seen and 5 unseen real-world tasks. From the obtained results in the figure, we have several key observations. 1) InSpire consistently enhances the success rates of the strong baseline VLA across 10 seen and 5 unseen tasks, achieving an average enhancement of 25% and 26%, respectively. 2) InSpire achieves notable performance gains and exhibits robustness against variations in object color, objects, and backgrounds in unseen tasks. Particularly, InSpire increases the relative success rate of the baseline by 100% on 4/5 of the unseen tasks. 3) InSpire incurs an additional time cost of 0.18 seconds per step on average compared to the baseline VLA. The additional computational overhead is primarily attributed to the spatial reasoning VQA task introducing more tokens into the model’s action generation process. However, considering the notable improvement in success rates, this trade-off is acceptable in a wide range of real-world robotic applications. 4) The backgrounds of the 15 real-world tasks are significantly more complex than those of the simulation tasks from the LIBERO environment, with numerous distracting elements like plastic bottles, baskets, and trash cans. Nevertheless, our InSpire demonstrates outstanding performance, consistent with the results observed in simulation experiments.
3.3 Comparison with State-of-the-art
In this part, we first apply InSpire to the miniVLA-VQ model and pretrain a 1B-parameter VLA on the LIBERO-90 dataset, which we refer to as InspireVLA-1B. We then compare our InspireVLA-1B with state-of-the-art reasoning-based VLAs, including SpatialVLA-4B[47]—a 4B-parameter VLA that injects 3D information into the input observations to achieve spatial-aware action prediction, and CoT-VLA-7B [48]—a 7B-parameter VLA that enhances action prediction through CoT reasoning.
Experimental Setup. We follow the common setup in [47, 48], leveraging tasks from LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long for evaluation. Each of SpatialVLA-4B, CoT-VLA-7B and our InspireVLA-1B is jointly fine-tuned on tasks from the four LIBERO datasets. We compare the performance on seen tasks from the four LIBERO datasets and directly borrow the results of SpatialVLA-4B and CoT-VLA-7B from their original papers. We perform LoRA fine-tuning [13] on InSpireVLA-1B using tasks from the four datasets, the hyperparameters adopted in our experiments are listed in Table 7. We report the average success rates and std errors on 3 random seeds. As in [48], we also compare InSpire with other three strong baselines: Diffusion Policy[9], Octo[26] and OpenVLA[15], their results are directly referred from [48] [^3] Since the LIBERO-pretrained checkpoints of Diffusion Policy, Octo, OpenVLA and CoT-VLA-7B in [48] have not been open-sourced, we are unable to compare the performance of those models on unseen tasks..
Table 2: Comparison with state-of-the-art. independently pretrained from scratch using the four LIBERO datasets. jointly finetuned on the four LIBERO datasets. The results of SpatialVLA-4B are obtained from its original paper [47], while the results of other rivals are referred from [48].
Model LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
@@ 3.3 Comparison with State-of-the-art
Results and Analysis. From the obtained results in Table 2, we have the following key observations. 1) Our InspireVLA-1B model consistently outperforms the five VLAs across the four LIBERO datasets. Particularly, InspireVLA-1B improves the two reasoning-based models SpatialVLA-4B and CoT-VLA-1B by 8.6% and 2.8%, with 4$\times$ and 7$\times$ parameter efficiency, respectively. 2) The three reasoning-based models (i.e., SpatialVLA-4B, CoT-VLA-7B and our InspireVLA-1B) achieve better performance than other three competitors, emphasizing the necessity of integrating an intermediate reasoning process into existing VLAs to capture the true causal relationships between observations and actions. 3) InspireVLA-1B demonstrates remarkable advantages on LIBERO-Spatial and LIBERO-Object, highlighting its effectiveness in transferring knowledge about spatial information and objects.
3.4 Ablations on Design Decisions
@@
Table 3: Ablation of the formulation of the spatial reasoning VQA task.
Setting             Question     Answer Seen Unseen
Baseline             -      - 83.8 3.6
1D Direct. $\mathsf{In\,which\,direction\,is\,the\,[object]\,relative\,to\,the\,robot?}$ $\mathsf{right/left/.../back/grasped}$ 90.8 18.0
3D Direct. $\mathsf{In\,which\,direction\,is\,[object]\,relative\,to\,the\,robot?\,x,y,z:}$ $\mathsf{[right,up,front]/.../grasped}$ 90.0 16.9
Proximity $\mathsf{What\,is\,the\,distance\,between\,the\,robot\,and\,[object]?}$ $\mathsf{far/middle/near/grasped}$ 88.7 9.3
3D Locat. $\mathsf{What\,is\,the\,accurate\,posi.\,of\,[object]\,relat.\,to\,the\,robot?\,x,y,z:}$ $\mathsf{[1,-3,4]/.../[2,0,1]}$ 88.9 10.1
Distance $\mathsf{What\,is\,the\,accurate\,distance\,between\,the\,robot\,and\,[object]?}$ $\mathsf{0/1/2/3/.../9}$ 82.8 6.0
Figure 5: Ablation of the insertion position of the spatial reasoning VQA task.
VQA Formulations. InSpire leverages a spatial reasoning VQA task as the bridging “language” between observations and actions. This implies that the formulation of the VQA task dictates the types of spatial information used by InSpire. Here, we investigate the impacts of different VQA formulations on InSpire’s performance. As shown in Table 3, we design five VQA tasks to introduce diverse spatial information, denoted as “1D Direction”, “3D Direction”, “Proximity”, “3D Location” and “Distance”. The performance of our InSpire approach integrated with each of these VQA tasks is reported in the table. We have the following observations from the results in the table. 1) Except for “Distance”, which achieves performance comparable to the baseline on seen tasks, the other four VQA formulations consistently outperform the baseline on both seen and unseen tasks. 2) “1D Direction” achieves the best performance among all VQA formulations, surpassing the baseline by large margins. 3) “3D Direction” demonstrates performance close to that of “1D Direction”, with both significantly outperforming the other VQA formulations. These results suggest that direction prediction-based VQA tasks are more effective in guiding the action generation process.
VQA Insertion Positions. In our developed InSpire approach, the tokens for the spatial reasoning VQA task are appended following the input visual observation tokens and preceding the language instruction tokens (Fig. 2). In this experiment, we investigate the impacts of different insertion positions of those VQA tokens on performance. Figure 5 presents a quantitative comparison of InSpire’s performance with two design decisions: inserting VQA tokens before the language instruction tokens (“VQA-First”) and placing them after the language instruction tokens (“Instrut-First”). As seen, “VQA-First” achieves superior performance on seen tasks compared to “Instrut-First”, but lags behind on unseen tasks. Moreover, InSpire substantially boosts the baseline model’s performance in both settings, highlighting the benefits of the spatial reasoning VQA task for enhancing spatial reasoning.
Figure 6: Qualitative Results. More qualitative results in both simulation and real-world environments are available at: https://koorye.github.io/proj/Inspire.
3.5 How Does InSpire Help Resolve Spurious Correlations?
@@
Addressing Shortcut Learning of VLAs. As shown in Task A, the baseline VLA, upon observing the “drawer” frequently seen during training, directly approaches it based on shortcut learning without understanding the input language instruction and visual information. In other words, the VLA learns spurious correlations between task-irrelevant features of observations and actions, leading to overfitting to seen tasks and poor generalization on unseen tasks. Our InSpire approach corrects the shortcut learning exhibited by the baseline, avoiding the neglect of essential elements like language instructions and spatial relations.
@@
Improving Model Robustness to Distractors. As illustrated in Task B, the baseline VLA struggles to identify task-specific objects in complex scenarios with various distractors, limiting its ability to produce correct actions. The primary reason may be that the direct observation-to-action mapping mechanism prevents the learned model from capturing causal relationships between observations and actions, making it sensitive to familiar distractors when predicting actions. By employing a spatial reasoning VQA task as the bridging “language” between observations and actions, our InSpire approach equips the VLA with spatial reasoning capability, thereby helping the VLA differentiate task-relevant objects from distractors.
@@
Enabling Continual Action Correction. Continual correction of incorrect actions is an essential skill for both humans and robots. Yet, as presented in Task C, the baseline VLA fails to correct the incorrect action and proceeds with following actions. This suggests the learned VLA uses perceptual data for decision-making in a very different way from how humans do. Conversely, our InSpire assists the model in continuously correcting errors until the task is completed. This is primarily because InSpire performs spatial reasoning for each visual observation along the trajectory, which enhances the model’s awareness of task execution status (e.g., the relative direction of the target object to the robot, the status of the gripper) and assists in correcting incorrect actions.
4 Conclusion, Limitations and Future Work
This work proposes Intrinsic Spatial Reasoning (InSpire) to mitigate the adverse effects of spurious correlations on the action prediction of VLAs. By incorporating a spatial reasoning VQA task as the bridging “language” between visual observations and low-level actions, InSpire endows VLAs with spatial reasoning capabilities without employing extra data or interacting with other large models. Notably, InSpire can be used as a plugin to improve existing autoregressive VLAs. Comprehensive experiments demonstrate InSpire’s effectiveness and flexibility, achieving consistent improvements on two state-of-the-art VLAs on both seen and unseen tasks in simulation and real-world environments. We hope this work provides a perspective on enhancing the generalization and robustness of VLAs for downstream tasks.
Although the results are encouraging, two key limitations remain in this work: 1) Although InSpire has potential for integration with any LLM-based VLAs, its effectiveness in enhancing diffusion policy-based VLAs, such as Octo [26] and $\pi_{0}$ [4], remains to be explored. Future work will focus on exploring strategies to facilitate seamless integration of InSpire into such models. 2) The relationships and differences in spurious correlations among different VLAs have not been investigated. VLAs pretrained with diverse paradigms and datasets may acquire varied spurious correlations. Conducting further empirical studies on these factors could improve the effectiveness and robustness of our approach in more complex robotic scenarios.
Appendix A Related Work
Vision-Language-Action Models. Vision-Language-Action models (VLAs), which are usually built upon these pretrained Vision-Language Models (VLMs) [14, 45, 36] have shown strong generalization capabilities in unseen objects, environments and instructions[6, 15, 33, 23]. A significant breakthrough came with RT-1 [5], which introduced a transformer-based framework for tokenizing visual inputs, language commands, and action outputs, achieving impressive zero-shot generalization through pretraining on 130k real-world robotic trajectories. This approach was subsequently advanced by RT-2 [6], which successfully transferred web-scale vision-language knowledge to robotic control domains. The performance boundaries are further pushed by RT-X [10], utilizing the expansive Open X-Embodiment dataset with 160k diverse robotic tasks to enhance model capabilities. OpenVLA [16] combines the powerful Llama 2 [36] model with a sophisticated visual processing system that integrates complementary features from both SigLIP [45] and DINO-v2 [31]. Octo [35] demonstrates remarkable efficiency, achieving fast adaptation to novel situations through a streamlined policy trained on the OXE dataset that runs effectively on consumer hardware. Another approach exemplified by $\pi_{0}$ [4] shares architectural similarities with RT-2 but distinguishes itself through enhanced cross-platform control capabilities derived from vision-language pretraining. $\pi_{0}$-FAST [32] accelerates training efficiency of $\pi_{0}$ by designing an efficient tokenizer for robotic actions and switching $\pi_{0}$’s diffusion-based prediction mode to autoregressive prediction.
Chain-of-Thought Reasoning. Chain-of-Thought (CoT) reasoning, initially pioneered in natural language processing [39], empowers models to decompose intricate problems into sequential, interpretable reasoning traces. Such techniques have recently been explored in the context of high-level task planning for robotic systems [19, 43, 27, 34, 28, 2]. Particularly in robotic manipulation, researchers have developed approaches using future state prediction through diffusion-based sub-goal planing [30], video generation [20, 48], while reinforcement learning systems incorporate reward shaping via CoT-guided exploration [46]. CoT-VLA [48] enhances VLAs with explicit visual CoT reasoning, where future image frames are autoregressively predicted as visual goals, followed by the generation of a short action sequence to achieve the planned goals. ECoT [42] conducts iterative reasoning about plans, sub-tasks, motions, and visually grounded elements such as object bounding boxes and end-effector positions before generating actions. Despite their progress, these methods often depend on complex intermediate computational processes or utilize external models for step-by-step reasoning, which can, to some extent, compromise their efficiency. RT-H [2] is the closest related work, guiding the policy to learn common low-level motion patterns across different tasks by using language motion prediction as an intermediate stage between tasks and actions. Unlike RT-H, our InSpire framework employs a spatial reasoning VQA task to explicitly associate objects of interest with corresponding actions, effectively reducing the impact of spurious correlations during learning.

### v3

@@ III-A1 LIBERO Evaluation
LIBERO-90 -Spatial -Object -Goal -Long Average
@@
We have the following key observations from Table I. 1) Our InSpire achieves the best or comparable performance compared to the two strong baseline VLAs on both seen and unseen tasks. Remarkably, InSpire enhances the absolute success rates of miniVLA-VQ by 6.2% and 10% on seen and unseen tasks, respectively. 2) The established performance gains of InSpire on $\pi_{0}$-FAST are not as significant as that on miniVLA-VQ. The primary reason could be that we train miniVLA-VQ from scratch but finetune the pretrained $\pi_{0}$-FAST. With substantially less data than $\pi_{0}$-FAST’s pretraining dataset, LIBERO-90 limits the ability of InSpire to enhance $\pi_{0}$-FAST’s spatial reasoning capability. 3) InSpire yields significantly better performance on LIBERO-Spatial and LIBERO-Object tasks compared to LIBERO-Goal and LIBERO-Long tasks, demonstrating its strength in handling spatial relationships and object interactions. 4) SpatialVLA, and other VLAs, exhibit poor generalization performance on tasks from the four unseen datasets, indicating these models’ inability to capture true causal relationships between observations and actions, leading to poor generalization beyond the distribution of their training data.
III-A2 CALVIN Evaluation
[…]
Quantitative results are shown in Fig.4. As observed, our InSpire consistently outperforms all baseline methods in completing 1 to 5 sequential tasks. This strong long-horizon capability is further highlighted by the average number of consecutively completed tasks, where InSpire again surpasses all competitors. This superior performance stems from our method’s effectiveness in mitigating spurious correlations, leading to more accurate and robust robot actions.
Figure 4: CALVIN Performance.
III-B Real-world Experiments
@@
TABLE II: Comparison with state-of-the-art. independently pretrained from scratch using the four LIBERO datasets. jointly finetuned on the four LIBERO datasets. The results of SpatialVLA-4B are obtained from its original paper [44], while the results of other rivals are referred from [45].
Model -Spatial -Object -Goal -Long Average
@@
Due to the difficulty in collecting huge pools of real-world behavioral demonstrations, it is challenging to train a model that displays generalization and robustness on varied real-world tasks, whether starting from scratch or finetuning a pretrained model lacking exposure to substantial real-world manipulation data. Hence, we employ $\pi_{0}$-FAST, the current best-performing open-source VLA, as the baseline model for real-world evaluations. We carefully design 10 seen tasks and 5 unseen tasks that focus on evaluating the VLA’s performance along multiple dimensions: spatial reasoning, interacting with novel objects and scenes, and following unknown instructions. The $\pi_{0}$-FAST model is full-finetuned using the 10 seen tasks, each consisting of training 10 training trajectories. During inference, we conduct 10 trials per task and randomize the configurations and orientations of task-specific objects for each trial. We use an AGILEX PiPER 6DOF robot arm. Other hyperparameters are consistent with those used in the simulation experiments.
Fig. 5 reports the success rates as well as time costs of $\pi_{0}$-FAST [32] integrated w/ or w/o our InSpire on 10 seen and 5 unseen real-world tasks. From the obtained results in the figure, we have several key observations. 1) InSpire consistently enhances the success rates of the strong baseline VLA across 10 seen and 5 unseen tasks, achieving an average enhancement of 25% and 26%, respectively. 2) InSpire achieves notable performance gains and exhibits robustness against variations in object color, objects, and backgrounds in unseen tasks. Particularly, InSpire increases the relative success rate of the baseline by 100% on 4/5 of the unseen tasks. 3) InSpire incurs an additional time cost of 0.18 seconds per step on average compared to the baseline VLA. The additional computational overhead is primarily attributed to the spatial reasoning VQA task introducing more tokens into the model’s action generation process. However, considering the notable improvement in success rates, this trade-off is acceptable in a wide range of real-world robotic applications. 4) The backgrounds of the 15 real-world tasks are significantly more complex than those of the simulation tasks from the LIBERO environment, with numerous distracting elements like plastic bottles, baskets, and trash cans. Nevertheless, our InSpire demonstrates outstanding performance, consistent with the results observed in simulation experiments.
TABLE III: Ablation of the formulation of the spatial reasoning VQA task.
Setting                      Question        Answer Seen Unseen
Baseline                        -          - 83.3±1.2 3.6±0.4
1D Direct. $\mathsf{In\,which\,direction\,is\,the\,[object]\,relative\,to\,the\,robot?}$ $\mathsf{right/left/.../back/grasped}$ 89.5±1.5 13.6±3.9
3D Direct. $\mathsf{In\,which\,direction\,is\,[object]\,relative\,to\,the\,robot?\,x,y,z:}$ $\mathsf{[right,up,front]/.../grasped}$ 87.9±2.1 15.0±1.9
Proximity $\mathsf{What\,is\,the\,distance\,between\,the\,robot\,and\,[object]?}$ $\mathsf{far/middle/near/grasped}$ 88.7±1.1 10.3±1.0
3D Locat. $\mathsf{What\,is\,the\,accurate\,posi.\,of\,[object]\,relat.\,to\,the\,robot?\,x,y,z:}$ $\mathsf{[1,-3,4]/.../[2,0,1]}$ 88.1±0.8 10.1±2.6
Distance $\mathsf{What\,is\,the\,accurate\,distance\,between\,the\,robot\,and\,[object]?}$ $\mathsf{0/1/2/3/.../9}$ 85.1±2.2 6.6±0.6
III-C Comparison with State-of-the-art
In this part, we first apply InSpire to the miniVLA-VQ model and pretrain a 1B-parameter VLA on the LIBERO-90 dataset, which we refer to as InspireVLA-1B. We then compare our InspireVLA-1B with state-of-the-art reasoning-based VLAs, including SpatialVLA-4B[44]—a 4B-parameter VLA that injects 3D information into the input observations to achieve spatial-aware action prediction, and CoT-VLA-7B [45]—a 7B-parameter VLA that enhances action prediction through CoT reasoning. We follow the common setup in [44, 45], leveraging tasks from LIBERO-Spatial, -Object, -Goal and -Long for evaluation. Each of SpatialVLA-4B, CoT-VLA-7B and our InspireVLA-1B is jointly fine-tuned on tasks from the four LIBERO datasets. We compare the performance on seen tasks from the four datasets and directly borrow the results of SpatialVLA-4B and CoT-VLA-7B from the original papers. We perform LoRA fine-tuning [13] on InSpireVLA-1B using tasks from the four datasets. We report the average success rates and std errors over 3 random seeds. As in [45], we also compare InSpire with other three strong baselines: Diffusion Policy[9], Octo[26] and OpenVLA[17], their results are directly referred from [45] [^1] Since the LIBERO-pretrained checkpoints of Diffusion Policy, Octo, OpenVLA and CoT-VLA-7B in [45] have not been open-sourced, we are unable to compare the performance of those models on unseen tasks..
[…]
III-D Ablations on Design Decisions
@@
Figure 6: Ablation study on the insertion position of the spatial reasoning VQA task.
III-D1 VQA Formulations
InSpire leverages a spatial reasoning VQA task as the bridging “language” between observations and actions. This implies that the formulation of the VQA task dictates the types of spatial information used by InSpire. Here, we investigate the impacts of different VQA formulations on InSpire’s performance. As shown in Table III, we design five VQA tasks to introduce diverse spatial information, denoted as “1D Direction”, “3D Direction”, “Proximity”, “3D Location” and “Distance”. The performance of our InSpire approach integrated with each of these VQA tasks is reported in the table. We have the following observations from the results in the table. 1) The four VQA formulations consistently outperform the baseline on both seen and unseen tasks. 2) “1D Direction” achieves the best performance among all VQA formulations, surpassing the baseline by large margins. 3) “3D Direction” demonstrates performance close to that of “1D Direction”, with both significantly outperforming the other VQA formulations. This suggest that direction prediction-based VQA tasks are more effective in guiding the action generation process.
III-D2 VQA Insertion Positions
In our developed InSpire approach, the tokens for the spatial reasoning VQA task are appended following the input vidsual observation tokens and preceding the language instruction tokens (Fig. 2). In this experiment, we investigate the impacts of different insertion positions of those VQA tokens on performance. Fig. 6 presents a quantitative comparison of InSpire’s performance with two design decisions: inserting VQA tokens before the language instruction tokens (“VQA-First”) and placing them after the language instruction tokens (“Instruct-First”). As seen, “VQA-First” achieves superior performance on seen tasks compared to “Instrut-First”, but lags behind on unseen tasks. Moreover, InSpire substantially boosts the baseline’s performance in both settings, highlighting the benefits of the auxiliary VQA task for enhancing spatial reasoning.
Figure 7: Qualitative Results. (Left) Action sequences and (Right) attention maps produced by the baseline model (mini-VLA-VQ [3]) integrated w/ or w/o our proposed InSpire approach. More qualitative results in both simulation and real-world environments are available at: https://koorye.github.io/proj/Inspire
III-E How Does InSpire Help Resolve Spurious Correlations?
@@ III-E1 Addressing Shortcut Learning of VLAs
As shown in Fig. 7 (Left) Task A, the baseline VLA, upon observing the “drawer” frequently seen during training, directly approaches it based on shortcut learning without understanding the input language instruction and visual information. In other words, the VLA learns spurious correlations between task-irrelevant features of observations and actions, leading to overfitting to seen tasks and poor generalization on unseen tasks. Our InSpire approach corrects the shortcut learning exhibited by the baseline, avoiding the neglect of essential elements like language instructions and spatial relations.
III-E2 Improving Model Robustness to Distractors
@@
As illustrated in Fig. 7 (Left) Task B, the baseline VLA struggles to identify task-specific objects in complex scenarios with various distractors, limiting its ability to produce correct actions. The primary reason may be that the direct observation-to-action mapping mechanism prevents the learned model from capturing causal relationships between observations and actions, making it sensitive to familiar distractors when predicting actions. By employing a spatial reasoning VQA task as the bridging “language” between observations and actions, InSpire equips the VLA with spatial reasoning capability, thereby helping the VLA differentiate task-relevant objects from distractors (as shown in Fig. 7 (Right)).
III-E3 Enabling Continual Action Correction
@@
Continual correction of incorrect actions is an essential skill for both humans and robots. Yet, as shown in Fig. 7 (Left) Task C, the baseline VLA fails to correct the incorrect action and proceeds with following actions. This suggests the learned VLA uses perceptual data for decision-making in a very different way from how humans do. Conversely, InSpire assists the model in continuously correcting errors until the task is completed. This is primarily because InSpire performs spatial reasoning for each visual observation along the trajectory, which enhances the model’s awareness of task execution status (e.g., the relative direction of the target object to the robot, the status of the gripper) and assists in correcting incorrect actions.
IV Related Work
IV-1 Vision-Language-Action Models (VLAs)
VLAs, built upon pretrained Vision-Language Models (VLMs) [15, 42], demonstrate strong generalization in robotics [6, 17]. This line of research, initiated by transformer-based frameworks like RT-1 [5] and advanced by transferring web-scale knowledge in RT-2 [6], has scaled to large datasets such as Open X-Embodiment [10]. Recent architectures leverage powerful foundation models (OpenVLA [18]), target computational efficiency (Octo [34]), or refine training and control strategies ($\pi_{0}$[4], $\pi_{0}$-FAST [32]).
IV-2 Chain-of-Thought (CoT) Reasoning
CoT reasoning has been adapted from NLP to high-level robotic task planning [40, 14, 33, 20, 30, 16]. Applications include sub-goal planning via diffusion or video generation [31] and reward shaping in reinforcement learning [43]. Models like CoT-VLA [45] and ECoT [39] use visual goal prediction and iterative reasoning, respectively, but can be inefficient by depending on complex intermediate computations or external models. The most related work, RT-H [2], uses language motion prediction as an intermediate step. In contrast, our InSpire framework employs a spatial reasoning VQA task to explicitly associate objects of interest with corresponding actions, reducing the impact of spurious correlations.
IV-3 Learning Spatial Reasoning